Suppose a consumer x and y has the following utility function U=x^0.4y^0.8 if Price of good x and y 10 and 20 and income constraint is 500$. Find the quantities of x and y maximize utility
1
Expert's answer
2020-06-15T11:13:41-0400

The utility function is:



"U=x^{0.4}y^{0.8}"

Utility will be maximized when:



"\dfrac{MU_x}{MU_y} = \dfrac{P_x}{P_y}"

From the utility function:



"MU_x =0.4 x^{-0.6}y^{0.8}\\[0.3cm]
MU_y = 0.8x^{0.4}y^{-0.2}"

The prices are:



"P_x = 10\\[0.3cm]
P_y = 20"

Therefore:



"\dfrac{0.4x^{-0.6}y^{0.8}}{0.8x^{0.4}y^{-0.2}} = \dfrac{10}{20}\\[0.3cm]
y = x ..............\text{Eqn 1}"

Suppose the consumer buys x units at Px = 10 and y units at Py. If his income is $500, the budget line is:



"10x + 20y = 500"

Substituting equation 1 into the budget line:



"10x + 20x = 500\\[0.3cm]
30x = 500\\[0.3cm]
\color{red}{x^* = \dfrac{50}{3}}"

Since "x = y" , then:



"\color{red}{y^* = \dfrac{50}{3}}"